OpenStudy (anonymous):

help please. a line passes through the point (8,9) and has a slope of 3/4 write the equation in slope intercept form for this line

OpenStudy (austinl):

We have the standard form of the line, \(y=mx+b\) where m is slope, so now we have \(\displaystyle y=\frac{3}{4}x+b\) Next you would need to plug in the point given, \((8,9)\rightarrow(x,y)\) Think you can solve it from there?

OpenStudy (anonymous):

okay, im still kind of confused I don't know what to put in for b

OpenStudy (anonymous):

would b be 9?